OpenStudy (anonymous):

3^2+2^2 Do i multiply the exponents or add them?

OpenStudy (ghazi):

neither add nor multiply

OpenStudy (unklerhaukus):

\[3^2+2^2=(3\times3)+(2\times2)\]

OpenStudy (cwrw238):

maybe you mean multiply?

OpenStudy (ghazi):

just solve the squares and add the terms

OpenStudy (anonymous):

3x^2+2x^2

OpenStudy (unklerhaukus):

\[3x^2+2x^2=(3+2)x^2\]

OpenStudy (anonymous):

so i dont do anything to the exponents just leave the x as squared?

OpenStudy (mathmate):

If the bases are different, like 3 and 2 in this case, you cannot add or subtract the exponents. Think of the bases as houses. If you go into the same building (base), you can go up and down the floors. If you are in different buildings (bases), you cannot go up one floor to the other building.

OpenStudy (anonymous):

nice visual

OpenStudy (anonymous):

thanks! :D

OpenStudy (anonymous):

or think what you would do if \(x\) was 10 \[3\times 10^2+2\times 10^2=300+200=500=5\times 10^2\]

OpenStudy (unklerhaukus):